OpenStudy (anonymous):

Your piggy bank has a total of 18 coins in it; some are dimes and some are nickels. If you have a total of $1.40, how many nickels do you have? 14 12 10 8

OpenStudy (anonymous):

8, you can do this by process of elimination, think that 10 dimes = 1.00 and then you will have eight other coins you have not specified and 8 nickels = .40 so the two combined are 1.40

OpenStudy (whpalmer4):

D is number of dimes, N is number of nickels. D+N = 18 (total number of coins) 10*D + 5*N = 140 (total number of cents) D = 18-N 10*(18-N) + 5*N = 140 180 - 10N + 5N = 140 Solve for N to get number of nickels. Number of dimes is 18 - number of nickels.

OpenStudy (whpalmer4):

Okay, Luis, your piggy bank has $1,239.85 in dimes and nickels. 16531 coins in all. Quickly, how many dimes, how many nickels?

OpenStudy (whpalmer4):

I'm illustrating the utility of understanding how to solve the problem when there's more than a trivial amount involved. No new knowledge required if you did it the way you denigrated, but your approach would be, um, problematic.